OpenStudy (anonymous):

whats the volume of the solid obtained when rotating the region under the graph of the function f(x)= x^(13/12)?? on the interval [1,3]??? need help

OpenStudy (anonymous):

creeg go back to your previous request to get the solution

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

vol = INT pi f(x)^2 dx between x = 1 and 3 = INT pi x^169/144 dx = pi [x^313/144 * 144/313] btw 1 and 3 = = pi [3^313/144 * 144/313 -1^313/144 * 144/313 ]

OpenStudy (anonymous):

aww its all squashed up - i copied it but its come out like that

OpenStudy (anonymous):

just scroll down on the left to findn ur previous post

OpenStudy (anonymous):

rotated about which axis?